OpenStudy (anonymous):

ِA map is in the scale of 1:25000. A lake on the map has an area of 5cm^2 .Calculate the area of the actual lake and give your answer in square meter in standard form.

OpenStudy (anonymous):

Let's first bring the area of the lake to real size in cm.'s 5*25,000=125,000 cm.^2

OpenStudy (anonymous):

Now, cm. to meter. 125,000/100=1250 meters^2

OpenStudy (anonymous):

ok but 1:25000, is this written in cube or square i mean is it perimeter area or volume?

OpenStudy (anonymous):

It's just the scale. The scale can't be written in a square or cube from. It is just shows the relation between the size of the lake in real life to the map.

OpenStudy (anonymous):

so 1 is the scale of lake from 25000 of map Am i right?

OpenStudy (anonymous):

0.00004 is the scale.

OpenStudy (anonymous):

can u explain me how

OpenStudy (anonymous):

1/25,000=0.00004

OpenStudy (anonymous):

i really dont understand

OpenStudy (anonymous):

@Studying?